# Constrained Discrete Diffusion

### v2

@@ 5 Experiments
To empirically demonstrate the advantage provided by CDD, we compare our method against similarly sized autoregressive and diffusion-based language models. Specifically, we benchmark our performance on constrained text generation against autoregressive baselines: GPT-2 and Llama 3.2 [44, 45], and discrete diffusion baselines MDLM and UDLM [7, 8]. We intentionally select models of similar size to, and up to one magnitude larger than, our base discrete diffusion model, noting that we expect the performance of both classes of models to scale similarly with size. As demonstrated by Schiff et al. [8], MDLM outperforms UDLM on the natural language tasks discussed in Sections 5.1 and 5.3. Consequently, we compare to both models in Section 5.2 and solely MDLM for the following natural language tasks. We use UDLM as the base discrete diffusion model for Section 5.2, and MDLM in Sections 5.1 and 5.3. For each application, CDD uses configurations as described in [7, 8] unless otherwise specified. Additional experiment details are available in Appendix  C.
@@ 5.1 Natural Language Toxicity Mitigation
While LLMs have demonstrated remarkable capabilities in generating human like text, a byproduct of data driven generation is the inadvertent creation of harmful, offensive and dangerous text. Despite undergoing alignment finetuning, current state-of-the-art models still frequently generate harmful outputs. While post-processing methods can be effective in moderating generations, they can be circumvented by adversarial or malicious prompts, resulting in toxic output generation. In this experiment, we generate outputs conditioned on prompts from the RealToxicityPrompts dataset [46]. We select prefixes with toxicity score above 0.5, testing how these models perform on toxic inputs, and perplexity below 30.0, avoiding incoherent generations. To model the constraints, GPT-Neo [47] is finetuned on the Jigsaw Toxic Comment Classification Challenge [^2] https://www.kaggle.com/c/jigsaw-toxic-comment- classification-challenge. We evaluate on 1,000 samples, and discuss ALM hyper-parameters in Appendix  D, and runtime in Appendix  E.
@@
[…]
@@
Figure 3: Results across different toxicity thresholds, PPL percentiles, LLM-Judge coherence, and coherence degradation levels. PPL is evaluated with GPT-2-XL, and coherence metrics are assessed by the LLM-Judge. Bold and underlined values denote the best and second-best, respectively.
@@
Among the baselines, GPT2 (124M) and Llama (1B) achieve the lowest perplexity and highest coherence score, however, they frequently produce toxicity content resulting in high violations. Although PPLM and FUDGE (124M) [26, 15] are able to decrease the violations for all toxicity thresholds, they are unable to consistently prevent toxic generations while suffering from an increase in perplexity and a decrease in coherence. MDLM (110M) exhibits higher perplexity than GPT2 and lower coherence, and similarly shares a non-negligible amount of toxicity constraint violations. In contrast, CDD provides perfect constraint satisfaction at each toxicity thresholds, maintaining comparable perplexity scores to its base diffusion language model and other baselines. Additionally, while constraint imposition degrades the coherence of the controllable generation baselines, CDD observes an increase in coherence as evaluated by the the LLM-Judge. These findings are important as they show how, using projection-based control strategies, it is possible to achieve complete toxicity constraint satisfaction at desired thresholds.
@@ 5.2 Molecular Generation
Next, we show the ability of CDD to impose constraints on a scientific task, specifically in the domain of molecular generation. In this experiment we generate SMILES strings [48], a linear representation of a molecule’s structure with a vocabulary consisting of a limited vocabulary of forty symbols and a grammar dictating molecular validity. Recent advances in generative modeling have enabled the design of molecular sequences by leveraging techniques from natural language processing. Despite their impressive ability to optimize for specific chemical properties, these models often generate molecules that fall short of practical requirements, either by producing compounds that are difficult to synthesize or by closely mimicking existing structures. Furthermore, generated molecules often fail to maximize qualitative properties such as drug-likeness (QED) [49] that are critical to the practicality of the generations and used as a qualitative metric for the experiments discussed in this section.
@@
To address this challenge, our framework incorporates two distinct constraint mechanisms: (1) synthetic feasibility, which ensures that generated molecules can be synthesized in a laboratory setting [50], and (2) novelty, which guarantees that the generated molecules are not already present in training datasets. For each constraint, we implement a dedicated projection operator, described in Appendix C, that imposes hard limits on the generation process, resulting in final outputs that adhere precisely to our desired standards of synthetic accessibility or novelty.
@@
Synthetic accessibility constraints. Synthetic accessibility is commonly assessed using black-box, non-differentiable functions. To enable gradient-based optimization, we train a surrogate model on the QM9 dataset [51], where training labels are derived from RDKit’s synthetic accessibility score [52]. This surrogate acts as our measurement function $g(\cdot)$ that approximates the original non-differentiable synthetic accessibility scorer, allowing construction of the projection operator depicted in Figure 4. While our molecular generation framework employs this surrogate model to optimize for synthetic accessibility during training, the actual assessment of accessibility violations is conducted by a separate, black-box external model. This external evaluation rigorously measures the degree to which generated molecules comply with the synthetic accessibility criteria, using a series of thresholds ($\tau=3.0$, $3.5$, $4.0$, and $4.5$). For the discrete diffusion baselines, we adapt the guidance mechanisms provided by [8] for QED by retraining on synthetic accessibility labels for Classifier-Based Guidance (CBG) and Classifier-Free Guidance (CFG). A similar adaptation is applied for the autoregressive baselines employing FUDGE guidance [15]. The results, summarized in Figure 5 (left), confirm that CDD achieves perfect compliance across all thresholds. This 0% violation rate is achieved, all while generating a competitive number of valid molecules and exhibiting the highest drug-likeness scores.
Figure 4: SA projection illustration.
@@
[]delimited-[][$\uparrow$][ ↑ ] []delimited-[][$\uparrow$][ ↑ ] []delimited-[][$\uparrow$][ ↑ ] $\tau=3.0$ $\tau=3.5$ $\tau=4.0$ $\tau=4.5$
@@
[…]
@@
ARD-CFG $\gamma=3$$\dagger$ 92M 79 0.60 91.61
MDLMD-CFG $\gamma=3$$\dagger$ 92M 96 0.60 69.82
UDLMD-CFG $\gamma=5$$\dagger$ 92M 64 0.62 93.69
CDDD-CFG $\gamma=5$ (Ours) 92M 251 0.60 0.00
@@
MDLMD-CBG $\gamma=3$$\dagger$ 92M 117 0.58 72.08
UDLMD-CBG $\gamma=10$$\dagger$ 92M 64 0.61 93.59
CDDD-CBG $\gamma=10$ (Ours) 92M 355 0.59 0.00
Figure 5: Left: Results for synthetic accessibility constrained molecule generation constraints. QED and constraint violations are reported for only valid molecules, and novel molecules must be valid and have no violation ($\tau\leq 3.0$). Right: Results for novelty projection with and without QED guidance. Violation represents percentage of valid, but not novel molecule generations. QED is reported for only novel molecules. Results denoted with are as reported by Schiff et al. [8]. Bold and underlined values mark the best and second-best, respectively.
@@
We compare agasint baselines with no QED guidance alongside CFG and CBG QED guidance [8] to evaluate the impact of our novelty constraint. The results are shown in the right side of Figure 5. Notice how the CBG setting yields a 203.4% increase in novel molecule generation, while the CFG setting shows a 161.4% increase. Even without guidance, the method still produces 48.1% more novel molecules, with only a minimal reduction in the QED score. These results are significant because they demonstrate that CDD can effectively generate novel molecules while maintaining a high QED score, which is crucial for many scientific tasks.
@@ 5.3 Instruction Following
Counting GPT-2 (zero-shot) 124M 19.2 ± 0.8 57.8 100
@@
Lexical GPT-2 (zero-shot) 124M 26.6 ± 4.2 28.1 97.6
@@
Finally, the last application focuses on satisfying user-specified symbolic constraints, a task that many autoregressive models struggle with. We evaluate CCD against such models using two types of constraints. First, we assess the ability of the tested models to accurately answer how many letters are in a word, such as ‘How many R’s are in Strawberry?’. We additionally test lexical constraints based of the COLLIE dataset [53], prompting the models to generate word-indexed constrained sentences, such as ‘Generate a sentence where the third-to-last word is mankind’. For each constraint, the projection operator to imposes hard constraints on the generation, pushing the samples back onto the constraint set.
@@
We include more evaluation and experiments in the Appendix. Specifically, Appendix C details the experimental setup, Appendix D evaluates CDD’s robustness through an extensive hyper-parameter ablation, and Appendix E reports runtime comparisons with the baselines.
@@ Acknowledgments
This work was partially supported by NSF grants CAREER-2401285 and RI-2334936 and UVA’s National Security Data & Policy Institute, ODNI Contracting Activity #2024-24070100001 Its view and conclusions reflect those of the authors only. The work has also been supported by LLNL under Contract DE-AC52- 07NA27344 and by the LLNL-LDRD Program under Project No. 24-ERD-010 and 24-SI-008. This manuscript has been co-authored by Lawrence Livermore National Security, LLC under Contract No. DE-AC52-07NA27344 with the U.S. Department of Energy. The United States Government retains, and the publisher, by accepting the article for publication, acknowledges that the United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes.
Appendix A Limitations
There are two key limitations that are currently faced in the development of constrained discrete models. The first regards the size of currently available discrete diffusion models. While current discrete diffusion models provide comparable performance in fluency and perplexity to similarly sized autoregressive models, including state-of-the-art models such as Llama 3.2 1B, their generative capabilities fall short of large-scale language models. Larger scale, closed-source discrete diffusion models have been developed and boasted impressive performance [33], but larger open-source models are yet to be provided to the research community. It is expected that the release of such larger models will further boost the performance of our proposed methods.
The second key limitation concerns an inherent byproduct of constraint imposition: increased computational overhead. While unconstrained discrete diffusion language models have been shown to provide significant speed-ups over autoregressive language models due to the parallelization of token generation [9, 7, 54], the inclusion of constraints within the generation process may limit this benefit. Nevertheless, this trade-off is often justified, as imposing constraints is essential in scenarios where correctness outweighs generation speed, such as ensuring safety-critical outputs or supporting rigorous scientific exploration. In such contexts, achieving accurate and reliable outcomes is prioritized over computational efficiency.
Appendix B Expanded Related Work
@@
A prominent strategy in controllable text generation is to employ constrained decoding algorithms, which incrementally parse outputs under specified formatting or structure requirements. Such approaches have been extensively used in tasks like code generation [55], semantic parsing [56], SQL queries [57], and grammar-constrained text generation [11]. These methods typically prune the vocabulary at each decoding step to only those tokens that comply with the target representation’s syntax or semantics. For instance, Geng et al. confine the logit distribution to tokens valid within a formal grammar, then employ a completion engine to extend partial sequences while maintaining grammatical consistency. Meanwhile, approaches that treat the original language model as a black box resort to external constraints. For example, [12] harnesses a secondary model to parse and restrict outputs, an option that becomes necessary when direct access to the logits is unavailable. However, these workarounds may suffer performance penalties due to their indirect control over the primary model’s probability distribution.
Beyond syntax-focused constraints, lexical constraints have also been extensively studied. Lu et al. [13] devise Neurologic decoding, which encourages the presence or absence of specific words in generated text via a modified beam search. They later improve adherence to these constraints by employing an A* search variant [14]. Similar frameworks combine toxicity classifiers with beam search to steer generation toward more benign text [26, 58]. Although these methods can effectively address a range of constraints, they generally rely on augmented sampling heuristics or additional classifiers rather than natively integrating constraints into the generation process.
@@
Another line of work has examined mapping natural language into logical representations to ensure factual and logical consistency [16, 17, 18]. By encoding linguistic inputs in a formal system, these methods can detect inconsistencies or factual errors and thereby provide a measure of control over generated outputs. For instance, Olausson et al., Jiang et al., and Zhang et al. convert natural language statements into first-order logic (FOL) and leverage theorem proving to verify factual correctness. However, FOL struggles to capture complex temporal or sequential dependencies, limiting its applicability to planning or process-oriented tasks. Similarly, other work uses entailment models to verify that a conclusion logically follows from a set of premises [62], though these methods generally stop short of actively generating text that satisfies domain-specific constraints.
@@
In robotics and embodied AI, natural language planning frequently incorporates specialized external modules to handle feasibility constraints. For example, Chen et al. [19] and Lin et al. [20] defer to domain-specific motion planners to generate viable trajectories, embedding constraint satisfaction directly into the underlying optimization routines. Although these solutions can integrate strong, domain-aware constraints, they do so by delegating control to highly specialized components. On the discrete side, Yang et al. [63] achieves constraint satisfaction by running a validation module at each step of plan construction and discarding infeasible actions through rejection sampling. While effective in certain settings, such methods typically require domain-specific expertise and do not uniformly integrate constraint enforcement into the generative model itself.
Appendix C Experimental Setup
C.1 Toxicity Mitigation

### v3

@@ 5 Experiments
To empirically demonstrate the advantage provided by CDD, we compare our method against similarly sized autoregressive and diffusion-based language models. Specifically, we benchmark our performance on constrained text generation against autoregressive baselines: GPT-2 and Llama 3.2 [radford2019language, grattafiori2024llama3herdmodels], and discrete diffusion baselines MDLM and UDLM [sahoo2024simple, schiff2024simple]. We intentionally select models of similar size to, and up to one magnitude larger than, our base discrete diffusion model, noting that we expect the performance of both classes of models to scale similarly with size. As demonstrated by schiff2024simple, MDLM outperforms UDLM on the natural language tasks discussed in Sections 5.1 and 5.3. Consequently, we compare to both models in Section 5.2 and solely MDLM for the following natural language tasks. We use UDLM as the base discrete diffusion model for Section 5.2, and MDLM in Sections 5.1 and 5.3. For each application, CDD uses configurations as described in [sahoo2024simple, schiff2024simple] unless otherwise specified. Additional experiment details are available in Appendix  B.
@@ 5.1 Natural Language Toxicity Mitigation
While LLMs have demonstrated remarkable capabilities in generating human like text, a byproduct of data driven generation is the inadvertent creation of harmful, offensive and dangerous text. Despite undergoing alignment finetuning, current state-of-the-art models still frequently generate harmful outputs. While post-processing methods can be effective in moderating generations, they can be circumvented by adversarial or malicious prompts, resulting in toxic output generation. In this experiment, we generate outputs conditioned on prompts from the RealToxicityPrompts dataset [gehman2020realtoxicityprompts]. We select prefixes with toxicity score above 0.5, testing how these models perform on toxic inputs, and perplexity below 30.0, avoiding incoherent generations. To model the constraints, GPT-Neo [black2021gpt] is finetuned on the Jigsaw Toxic Comment Classification Challenge [^2] https://www.kaggle.com/c/jigsaw-toxic-comment- classification-challenge. We evaluate on 1,000 samples, and discuss ALM hyper-parameters in Appendix  C, runtime in Appendix  D, and diversity as measured by entropy [zheng2024masked] in Appendix  E.
@@
$\text{GPT2}_{\text{PPLM}}$ 124M 46.40 34.05 18.88 -55.8 16.1 8.4 4.0
$\text{GPT2}_{\text{FUDGE}_{\lambda=2}}$ 124M 26.46 18.79 41.79 -2.1 31.5 19.7 12.7
$\text{GPT2}_{\text{FUDGE}_{\lambda=9}}$ 124M 81.84 19.22 38.90 -8.9 30.6 19.6 11.7
@@
Figure 3: Results across different toxicity thresholds, PPL percentiles, LLM‑Judge coherence, and coherence degradation levels. PPL is evaluated with GPT‑2‑XL, and coherence metrics are assessed by the LLM‑Judge. Bold and underlined values denote the best and second‑best, respectively.
@@
Among the baselines, GPT2 (124M) and Llama (1B) achieve the lowest perplexity and highest coherence score, however, they frequently produce toxicity content resulting in high violations. Although PPLM and FUDGE (124M) [dathathri2019plug, Yang_2021] are able to decrease the violations for all toxicity thresholds, they are unable to consistently prevent toxic generations while suffering from an increase in perplexity and a decrease in coherence. MDLM (110M) exhibits higher perplexity than GPT2 and lower coherence, and similarly shares a non-negligible amount of toxicity constraint violations. In contrast, CDD provides perfect constraint satisfaction at each toxicity thresholds, maintaining comparable perplexity scores to its base diffusion language model and other baselines. Additionally, while constraint imposition degrades the coherence of the controllable generation baselines, CDD observes an increase in coherence as evaluated by the LLM-Judge. These findings are important as they show how, using projection-based control strategies, it is possible to achieve complete toxicity constraint satisfaction at desired thresholds.
@@ 5.2 Molecular Generation
Next, we show the ability of CDD to impose constraints on a scientific task, specifically in the domain of molecular generation. In this experiment we generate SMILES strings [weininger1988smiles], a linear representation of a molecule’s structure with a vocabulary consisting of a limited vocabulary of forty symbols and a grammar dictating molecular validity. Recent advances in generative modeling have enabled the design of molecular sequences by leveraging techniques from natural language processing. Despite their impressive ability to optimize for specific chemical properties, these models often generate molecules that fall short of practical requirements, either by producing compounds that are difficult to synthesize or by closely mimicking existing structures. Furthermore, generated molecules often fail to maximize qualitative properties such as drug-likeness (QED) [bickerton2012quantifying] that are critical to the practicality of the generations and used as a qualitative metric for the experiments discussed in this section.
@@
To address this challenge, our framework incorporates two distinct constraint mechanisms: (1) synthetic feasibility, which ensures that generated molecules can be synthesized in a laboratory setting [ertl2009estimation], and (2) novelty, which guarantees that the generated molecules are not already present in training datasets. For each constraint, we implement a dedicated projection operator, described in B, that imposes hard limits on the generation process, resulting in final outputs that adhere precisely to our desired standards of synthetic accessibility or novelty.
@@
Synthetic accessibility constraints. Synthetic accessibility is commonly assessed using black-box, non-differentiable functions. To enable gradient-based optimization, we train a surrogate model on the QM9 dataset [ruddigkeit2012enumeration], where training labels are derived from RDKit’s synthetic accessibility score [bento2020open]. This allows us to approximate a differentiable function $g(\cdot)$ for synthetic accessibility. While our molecular generation framework employs this surrogate model to optimize for synthetic accessibility during training, the actual assessment of accessibility violations is conducted by a separate, black-box external model. This external evaluation rigorously measures the degree to which generated molecules comply with the synthetic accessibility criteria, using a series of thresholds ($\tau=3.0$, $3.5$, $4.0$, and $4.5$). For the discrete diffusion baselines, we adapt the guidance mechanisms provided by [schiff2024simple] for QED by retraining on synthetic accessibility labels for Classifier-Based Guidance (CBG) and Classifier-Free Guidance (CFG). A similar adaptation is applied for the autoregressive baselines employing FUDGE guidance [Yang_2021]. The results, summarized in Figure 4 (left), confirm that CDD achieves perfect compliance across all thresholds. This 0% violation rate is achieved, all while generating a competitive number of valid molecules and exhibiting the highest drug-likeness scores.
@@
$\tau=3.0$ $\tau=3.5$ $\tau=4.0$ $\tau=4.5$
@@
$\textbf{CDD}_{\tau=3.0}\textbf{(Ours)}$ 353 36 0.63 0.0 0.0 0.0 0.0
$\textbf{CDD}_{\tau=3.5}\textbf{(Ours)}$ 863 22 0.62 0.0 0.0 0.0
$\textbf{CDD}_{\tau=4.0}\textbf{(Ours)}$ 936 31 0.61 0.0 0.0
$\textbf{CDD}_{\tau=4.5}\textbf{(Ours)}$ 938 33 0.58 0.0
@@
ARD‑CFG $\gamma=3\dagger$ 92M 79 0.60 91.61
MDLMD‑CFG $\gamma=3\dagger$ 92M 96 0.60 69.82
UDLMD‑CFG $\gamma=5\dagger$ 92M 64 0.62 93.69
CDDD‑CFG $\gamma=5$ (Ours) 92M 251 0.60 0.00
@@
MDLMD‑CBG $\gamma=3\dagger$ 92M 117 0.58 72.08
UDLMD‑CBG $\gamma=10\dagger$ 92M 64 0.61 93.59
CDDD‑CBG $\gamma=10$ (Ours) 92M 355 0.59 0.00
Figure 4: Left: Results for synthetic accessibility constrained molecule generation constraints. QED and constraint violations are reported for only valid molecules, and novel molecules must be valid and have no violation ($\tau\leq 3.0$). Right: Results for novelty projection with and without QED guidance. Violation represents percentage of valid, but not novel molecule generations. QED is reported for only novel molecules. Results denoted with are as reported by schiff2024simple. Bold and underlined values mark the best and second-best, respectively.
@@
We compare against baselines with no QED guidance alongside CFG and CBG QED guidance [schiff2024simple] to evaluate the impact of our novelty constraint. The results are shown in the right side of Figure 4. Notice how the CBG setting yields a 203.4% increase in novel molecule generation, while the CFG setting shows a 161.4% increase. Even without guidance, the method still produces 48.1% more novel molecules, with only a minimal reduction in the QED score. These results are significant because they demonstrate that CDD can effectively generate novel molecules while maintaining a high QED score, which is crucial for many scientific tasks.
@@ 5.3 Instruction Following
Counting
GPT-2 (zero-shot) 124M 19.2 ± 0.8 57.8 100
@@
Lexical
GPT-2 (zero-shot) 124M 26.6 ± 4.2 28.1 97.6
@@
Finally, the last application focuses on satisfying user-specified symbolic constraints, a task that many autoregressive models struggle with. We evaluate CDD against such models using two types of constraints. First, we assess the ability of the tested models to accurately answer how many letters are in a word, such as ‘How many R’s are in Strawberry?’. We additionally test lexical constraints based of the COLLIE dataset [yao2023collie], prompting the models to generate word-indexed constrained sentences, such as ‘Generate a sentence where the third-to-last word is mankind’. For each constraint, the projection operator to imposes hard constraints on the generation, pushing the samples back onto the constraint set.
@@
We include more evaluation and experiments in the Appendix. Specifically, Appendix B details the experimental setup, Appendix C evaluates CDD’s robustness through an extensive hyper‑parameter ablation, and Appendix D reports runtime comparisons with the baselines.
6 Limitations
There are two key limitations that are currently faced in the development of constrained discrete models. The first regards the size of currently available discrete diffusion models. While current discrete diffusion models provide comparable performance in fluency and perplexity to similarly sized autoregressive models, including state-of-the-art models such as Llama 3.2 1B, their generative capabilities fall short of large-scale language models. Larger scale, closed-source discrete diffusion models have been developed and boasted impressive performance [shi2024simplified], but larger open-source models are yet to be provided to the research community. It is expected that the release of such larger models will further boost the performance of our proposed methods.
The second key limitation concerns an inherent byproduct of constraint imposition: increased computational overhead. While unconstrained discrete diffusion language models have been shown to provide significant speed-ups over autoregressive language models due to the parallelization of token generation [loudiscrete, sahoo2024simple, christopher2024speculative], the inclusion of constraints within the generation process may limit this benefit. Nevertheless, this trade-off is often justified, as imposing constraints is essential in scenarios where correctness outweighs generation speed, such as ensuring safety-critical outputs or supporting rigorous scientific exploration. In such contexts, achieving accurate and reliable outcomes is prioritized over computational efficiency.
@@ Acknowledgments
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No 2234693. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation. This work was partially supported by NSF grants CAREER-2401285, RI-2533631, RI-2334936, DARPA under Contract No. #HR0011252E005, and UVA’s National Security Data & Policy Institute, ODNI Contracting Activity #2024-24070100001 The work has also been supported by LLNL under Contract DE-AC52- 07NA27344 and by the LLNL-LDRD Program under Project No. 24-ERD-010 and 24-SI-008. This manuscript has been co-authored by Lawrence Livermore National Security, LLC under Contract No. DE-AC52-07NA27344 with the U.S. Department of Energy. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the NSF, DOE, DARPA, or DOD.
Appendix A Expanded Related Work
@@
A prominent strategy in controllable text generation is to employ constrained decoding algorithms, which incrementally parse outputs under specified formatting or structure requirements. Such approaches have been extensively used in tasks like code generation [poesia2022synchromesh], semantic parsing [shin-etal-2021-constrained], SQL queries [scholak-etal-2021-picard], and grammar-constrained text generation [geng-etal-2023-grammar]. These methods typically prune the vocabulary at each decoding step to only those tokens that comply with the target representation’s syntax or semantics. For instance, geng-etal-2023-grammar confine the logit distribution to tokens valid within a formal grammar, then employ a completion engine to extend partial sequences while maintaining grammatical consistency. Meanwhile, approaches that treat the original language model as a black box resort to external constraints. For example, [geng2024sketch] harnesses a secondary model to parse and restrict outputs, an option that becomes necessary when direct access to the logits is unavailable. However, these workarounds may suffer performance penalties due to their indirect control over the primary model’s probability distribution.
Beyond syntax-focused constraints, lexical constraints have also been extensively studied. lu2020neurologic devise Neurologic decoding, which encourages the presence or absence of specific words in generated text via a modified beam search. They later improve adherence to these constraints by employing an A* search variant [lu2021neurologic]. Similar frameworks combine toxicity classifiers with beam search to steer generation toward more benign text [dathathri2019plug, hartvigsen2022toxigen]. Although these methods can effectively address a range of constraints, they generally rely on augmented sampling heuristics or additional classifiers rather than natively integrating constraints into the generation process.
@@
Another line of work has examined mapping natural language into logical representations to ensure factual and logical consistency [gopalan2018sequence, liu2023grounding, liu2024logic]. By encoding linguistic inputs in a formal system, these methods can detect inconsistencies or factual errors and thereby provide a measure of control over generated outputs. For instance, olausson2023linc, jiang2024leanreasoner, and zhang2023improved convert natural language statements into first-order logic (FOL) and leverage theorem proving to verify factual correctness. However, FOL struggles to capture complex temporal or sequential dependencies, limiting its applicability to planning or process-oriented tasks. Similarly, other work uses entailment models to verify that a conclusion logically follows from a set of premises [weir2023nellie], though these methods generally stop short of actively generating text that satisfies domain-specific constraints.
@@
In robotics and embodied AI, natural language planning frequently incorporates specialized external modules to handle feasibility constraints. For example, chen2023autotamp and lin2023text2motion defer to domain-specific motion planners to generate viable trajectories, embedding constraint satisfaction directly into the underlying optimization routines. Although these solutions can integrate strong, domain-aware constraints, they do so by delegating control to highly specialized components. On the discrete side, yang2023plug achieves constraint satisfaction by running a validation module at each step of plan construction and discarding infeasible actions through rejection sampling. While effective in certain settings, such methods typically require domain-specific expertise and do not uniformly integrate constraint enforcement into the generative model itself.
Appendix B Experimental Setup
B.1 Toxicity Mitigation